OpenStudy (anonymous):

given 4(2^n-1)=16^n. find the value of n.

OpenStudy (anonymous):

Do you know how to do it?

OpenStudy (solomonzelman):

do the right side. 16^n= (4^2)^n= 4^(2n).

OpenStudy (solomonzelman):

now can you solve it?

OpenStudy (triciaal):